# Re-HOLD: Video Hand Object Interaction Reenactment via adaptive Layout-instructed Diffusion Model

### v2

@@ Abstract
Current digital human studies focusing on lip-syncing and body movement are no longer sufficient to meet the growing industrial demand, while human video generation techniques that support interacting with real-world environments (e.g., objects) have not been well investigated. Despite human hand synthesis already being an intricate problem, generating objects in contact with hands and their interactions presents an even more challenging task, especially when the objects exhibit obvious variations in size and shape. To cope with these issues, we present a novel video Reenactment framework focusing on Human-Object Interaction (HOI) via an adaptive Layout-instructed Diffusion model (Re-HOLD). Our key insight is to employ specialized layout representation for hands and objects, respectively. Such representations enable effective disentanglement of hand modeling and object adaptation to diverse motion sequences. To further improve the generation quality of HOI, we have designed an interactive textural enhancement module for both hands and objects by introducing two independent memory banks. We also propose a layout-adjusting strategy for the cross-object reenactment scenario to adaptively adjust unreasonable layouts caused by diverse object sizes during inference. Comprehensive qualitative and quantitative evaluations demonstrate that our proposed framework significantly outperforms existing methods. Project page: https://fyycs.github.io/Re-HOLD.
@@ 2 Related Work
Figure 2: Overview of our proposed Re-HOLD framework. We propose a two-branch framework that consists of a Reference U-Net and a Denoising U-Net. The Reference U-Net takes a reference object image for object texture encoding while the denoising one takes noise latent and layout guidance as input for diffusion processing. To enhance the quality of HOI generation, we adopt the HOI Restoration Module and a hand memory bank for hand information restoration. The object memory bank is designed to store fine-grained object information.
@@ 2.1 Human Body Animation
Recent studies have leveraged powerful diffusion-based models to address the challenge of human body animation. One prominent approach involves using a UNet-based network [42] enhanced with cross-attention mechanisms [46] to inject additional information. PIDM [1] was the first to introduce classifier-free diffusion guidance for pose-guided human image generation, as further developed by [32]. DreamPose [22] utilizes UV maps as motion signals and performs conditional embeddings to achieve motion transfer. Similar concepts have been explored in other works [20, 50, 56, 73, 4, 12, 58]. RealisDance [72] takes DWPose [60], SMPL [2], and HaMeR [36] as input to generate realistic hand poses. While these approaches have yielded promising results, there are challenges when dealing with accurate HOI generations. Recently, InteractDiffusion [18] introduces the layout input to capture intricate interaction relationships. Inspired by this, we employ the sparse layout guidance as the driving signal and adaptively modify the hand-object interaction information during inference.
@@ 2.2 Video Generation and Editing
Numerous studies have focused on adapting pre-trained diffusion models for video generation and editing, including zero-shot [9, 59, 3, 39, 61, 29] and one-shot-tuned [54, 74, 27, 11] learning frameworks. However, these approaches typically demand extensive per-video fine-tuning, which significantly hampers their practical application. Alternatively, other research [7, 5, 31, 52, 64, 37, 28] takes a training-based approach, where models are trained on large datasets to ensure they can serve as efficient editing tools during inference. However, none of these methods specifically address the adaptive generation of human-object interaction, and accurate HOI synthesis has always been a significant challenge in video generation tasks. In addition, a significant proportion of video generation and editing methodologies [40, 9, 10, 59, 39, 3, 61, 54, 27, 8, 43, 33, 44, 30] primarily depend on textual input for guidance during the editing process. However, textual prompts can sometimes fall short in accurately conveying the user’s intentions. For our specific application, utilizing an object image as a form of guidance proves to be more precise and effective.

### v3

@@ Abstract
Current digital human studies focusing on lip-syncing and body movement are no longer sufficient to meet the growing industrial demand, while human video generation techniques that support interacting with real-world environments (e.g., objects) have not been well investigated. Despite human hand synthesis already being an intricate problem, generating objects in contact with hands and their interactions presents an even more challenging task, especially when the objects exhibit obvious variations in size and shape. To tackle these issues, we present a novel video reenactment framework focusing on Human-Object Interaction (HOI) via an adaptive Layout-instructed Diffusion model (Re-HOLD). Our key insight is to employ specialized layout representation for hands and objects, respectively. Such representations enable effective disentanglement of hand modeling and object adaptation to diverse motion sequences. To further improve the quality of the HOI generation, we design an interactive textural enhancement module for both hands and objects by introducing two independent memory banks. We also propose a layout adjustment strategy for the cross-object reenactment scenario to adaptively adjust unreasonable layouts caused by diverse object sizes during inference. Comprehensive qualitative and quantitative evaluations demonstrate that our proposed framework significantly outperforms existing methods. Project page: https://fyycs.github.io/Re-HOLD.
@@ 2 Related Work
Figure 2: Overview of our proposed Re-HOLD framework. We propose a two-branch framework that consists of a Reference U-Net and a Denoising U-Net. The Reference U-Net takes a reference object image for object texture encoding while the denoising one takes noise latent and layout guidance as input for diffusion processing. To enhance the quality of HOI generation, we adopt the HOI Restoration Module for hand information and fine-grained object information restoration.
@@ 2.1 Human Body Animation
Recent studies have leveraged powerful diffusion-based models to address the challenge of human body animation. One prominent approach involves using a UNet-based network [44] enhanced with cross-attention mechanisms [48] to inject additional information. PIDM [1] was the first to introduce classifier-free diffusion guidance for pose-guided human image generation, as further developed by [33]. DreamPose [23] utilizes UV maps as motion signals and performs conditional embeddings to achieve motion transfer. Similar concepts have been explored in other works [21, 53, 59, 76, 4, 13, 61, 8]. RealisDance [75] takes DWPose [63], SMPL [2], and HaMeR [38] as input to generate realistic hand poses. While these approaches have yielded promising results, there are challenges when dealing with accurate HOI generations. Recently, InteractDiffusion [19] introduces the layout input to capture intricate interaction relationships. Inspired by this, we employ the sparse layout guidance as the driving signal and adaptively modify the hand-object interaction information during inference.
@@ 2.2 Video Generation and Editing
Numerous studies have focused on adapting pre-trained diffusion models for video generation and editing, including zero-shot [10, 62, 3, 41, 64, 30, 52] and one-shot-tuned [57, 78, 28, 12] learning frameworks. However, these approaches typically demand extensive per-video fine-tuning, which significantly hampers their practical application. Alternatively, other research [7, 5, 32, 55, 67, 39, 29, 77, 35] takes a training-based approach, where models are trained on large datasets to ensure they can serve as efficient editing tools during inference. However, none of these methods specifically address the adaptive generation of human-object interaction, and accurate HOI synthesis has always been a significant challenge in video generation tasks. In addition, a significant proportion of video generation and editing methodologies [42, 10, 11, 62, 41, 3, 64, 57, 28, 9, 45, 34, 46, 31] primarily depend on textual input for guidance during the editing process. However, textual prompts can sometimes fall short in accurately conveying the user’s intentions. For our specific application, utilizing an object image as a form of guidance proves to be more precise and effective.